OpenStudy (anonymous):

a wheel makes 3 7over8 revoltuions per minute. if ti rotates for 22 min, how many revolutions does it make?

OpenStudy (whpalmer4):

Is that \(3 \frac{7}{8}\) revolutions per minute? I'll assume it is. The procedure is the same, in any case. We have a rate problem: total revolutions = rate * time rate = \(3 \frac{7}8\) revolutions/minute time = 22 minutes total revolutions = \(3 \frac{7}8 * 22\) with a mixed fraction, convert it to an improper fraction by multiplying the whole number * denominator and adding the numerator — this gives you the numerator of the improper fraction. The denominator remains the same. \[3 \frac{7}8 \rightarrow \frac{3*8+7}{8} = \frac{31}{8}\]Now you can do the multiplication. Alternatively, convert the rate to a decimal and multiply.